# Deliberate Planning of 3D Bin Packing on Packing Configuration Trees

### v3

@@ 3.2 Markov Decision Process Formulation
The transition $\mathcal{P}(s_{t+1}|s_{t})$ is jointly determined by the current policy $\pi$ and the probability distribution of sampling items. Our online sequences are generated on the fly from an item set $\mathcal{I}$ in a uniform distribution. The transferability of our method on item sampling distributions different from the training one is discussed in Appendix C.
@@
Our reward function $R$ is defined as $r_{t}=c_{r}\cdot w_{t}$ once $n_{t}$ is inserted into PCT as an internal node successfully; otherwise, $r_{t}=0$ and the packing episode ends. Here, $c_{r}$ is a constant and $w_{t}$ is the weight of $n_{t}$. The choice of $w_{t}$ which depends on the customized needs. For simplicity and clarity, unless otherwise noted, we set $w_{t}$ as the volume $v_{t}$ of $n_{t}$, where $v_{t}=s^{x}_{n}\cdot s^{y}_{n}\cdot s^{z}_{n}$.
@@ 3.3 Recursive Packing for Large-Scale BPP
Figure 3: Problem decomposition based on sliding windows, illustrated in the $xoz$ plane. Low-resolution decomposition (a) results in the loss of solutions, while fine-grained partitioning (b) clearly increases computational overhead.
For problem decomposition, an intuitive approach is to divide the bin $C$ into uniform sections with resolution $r$ along each dimension, and maintain a sliding window that traverses the entire bin space in a convolution-like manner, resulting in smaller sub-bins $\mathbf{c}=\{c^{1},...,c^{n}\}$. Each $c^{i}$, along with its overlapped packed items $\mathbf{B}^{i}$ and empty space $\mathbf{L}^{i}$, is treated as a sub-problem $\mathcal{T}^{i}$. While this decomposition is intuitive, it’s not aware of the in-bin spatial distribution, leading to potential degradation in solution quality. Figure 3 (a) gives a demonstration for this, solutions that exist in the original bin $C$ are no longer feasible in any of the decomposed sub-bin $c^{i}$. Although finer-grained partitioning can to some extent avoid this (Figure 3 (b)), it also leads to a $O(r^{6})$ space complexity, significantly increasing the number of sub-problems and computational costs. Even more, for problems requiring decisions in continuous domains, fine-grained partitioning is inherently infeasible.
@@
Given sub-problems $\mathbf{T}$, the solution for $\mathcal{T}_{i}\in\mathbf{T}$ can be generated using a pre-trained policy $\pi_{\theta}(\cdot|\hat{\mathcal{T}_{i}},\hat{n})$. These solutions need to be integrated to conduct a global placement. Zhao et al. (2021) propose evaluating placement quality in multi-bin decision scenarios utilizing the learned state value approximator $V(c_{t},n_{t})=\mathbb{E}\left[\sum_{t}^{\infty}\gamma^{k}r_{t+k}\right]$, which captures the cumulative space utilization achievable within bin $c$ in the future after placing the current item $n$ in $c$. Here $\gamma\to 1$ is the reward discount factor. However, such a multi-bin decision approach assumes that bins are independent and do not affect each other, which obviously no longer holds in recursive packing. Moreover, the sub-bin only provides local placement evaluation, which does not necessarily represent the global optimality. As illustrated in Figure 5 (a), from the view of sub-bin $c_{1}$, the current item is compactly placed. However, from the global view of bin $C$ (Figure 5 (b)), unused space exists beneath the current item which cannot be utilized for subsequent packing due to the top-down robot packing requirement, indicating a sub-optimal solution.
Figure 5: From the view of sub-bin $c_{1}$ (a), the current item is optimally placed, but for the global view of bin $C$ (b), unused space beneath the item exists which can no longer be utilized.
@@
$l^{*}=\arg\max_{l\in\textbf{L}}\min_{c_{i}\in\mathbf{c}}\Phi(l,c_{i}).\vspace{-4pt}$ (9)
@@ 3.4 Uniform Planning for BPP Variations
Figure 7: Mainstream BPP variations–offline (a), buffering (b), and lookahead (c)–can be modeled as search trees without conflicting planning constraints. We solve them out of the box via a unified framework (d) powered by a pre-trained PCT model.
@@
We first formalize offline 3D-BPP, where all items are selectable, as a planning problem. An intuitive approach is performing a traversal tree search over all packing orders and positions for all items $\mathcal{I}$. Each path of the search tree represents a possible solution, and we choose the one with the highest accumulated space utilization $\sum_{i=0}^{|\mathcal{I}|}v_{i}$ for execution. This brute force search has a computational complexity of $O(|\mathcal{I}|!\cdot|\mathcal{A}|^{|\mathcal{I}|})$, where $|\mathcal{A}|$ represents action space size. Leveraging the pre-trained policy model $\pi_{\theta}$, which determines item position, the search can be simplified to only consider item order, as exhibited in Figure 7 (a), lowering the complexity to $O(|\mathcal{I}|!)$. The constraints for planning selectable items are as follows:
@@
    Items are placed based on the node sequence of the planned path. The planning only conducts once.
Buffering packing can be considered a direct extension of offline packing, where unknown items should be additionally considered. These items cannot be explicitly included as tree nodes. We use the state value function $V(\cdot)$ to implicitly estimate their distribution and future values. For $s$ items within $Fov_{c}$, their placement order can still be enumerated during the planning, while items outside of $Fov_{c}$ are all modeled as a single leaf node in the path end, as shown in Figure 7 (b). The value of each path is calculated by the sum of item volumes in the path and the value of the leaf node, i.e., $\sum_{i=0}^{s}v_{i}+V(\cdot)$. Since future arrival exist, the planning follows the principle of MBP, where only the first node of the selected path is executed and the search reinitializes when a new item arrives. The additional planning constraints when introducing unknown items are summarized as:
@@
We can find that these planning constraint for different operation attributes are compatible and can be integrated into a unified framework, which can be applied to general packing scenarios where $p\geq 0$, $s\geq 1$, and $u\geq 0$, as illustrated in Figure 7 (d). We denote this unified search tree as Tree of Packing (ToP). To ensure that the planning meets the real-time requirements of industrial packing, we introduce Monte Carlo Tree Search (MCTS) (Silver et al., 2016, 2017) which reduces the time complexity of the brute-force search from $O(|p+s|!)$ to $O((p+s)\cdot m)$. Here $m$ is the number of sampled paths. During planning, MCTS at adjacent time steps may share the same part of paths (i.e., item sequences). We maintain a global cache that stores previously visited paths to avoid redundant computations. This approach nearly halve the decision time costs.
@@ 4 Experiments
In this section, we first present PCT performance combined with different leaf node expansion schemes. We then highlight the advantages of the structured packing representation, including improved node spatial relation representations and a more flexible action space. Next, we validate the effectiveness of PCT-driven planners in solving industrial packing problems, specifically large-scale packing and different variations of BPP setting. Finally, we introduce our real-world packing robot in an industrial warehouse, carefully designed to meet constrained placement and transportation stability.
@@
Although there are very few online packing implementations publicly available, we still do our best to collect or reproduce various online 3D-BPP algorithms, both heuristic and learning-based, from potentially relevant literature. We help the heuristics to make pre-judgments of placement constraints, e.g., stability, in case of premature downtime. The learning-based agents are trained until there are no significant performance gains. All methods are implemented in Python and tested on 2000 instances with a desktop computer equipped with a Gold 5117 CPU and a GeForce TITAN V GPU.
@@
Some baselines (Karabulut and Inceoglu, 2004; Wang and Hauser, 2019b) need to traverse the entire coordinate space to find the optimal solution, and the running costs explode as the spatial discretization accuracy increases. To ensure that all algorithms are runnable within a reasonable period, we use the discrete dataset proposed by Zhao et al. (2021) without special declaration. The bin sizes $S^{d}$ are set to 10 with $d\in\{x,y,z\}$ and the item sizes $s^{d}\in\mathbb{Z}^{+}$ are not greater than $S^{d}/2$ to avoid over-simplification. Our performance on the continuous dataset will be reported in Section 4.3. Considering that there are many practical scenarios of 3D-BPP, we choose three representative ones:
[…]
@@ 4.1 Performance of PCT Policies
We first report the performance of PCT combined with different leaf node expansion schemes. Three existing schemes which have proven to be both efficient and effective are adopted here: Corner Point (CP), Extreme Point (EP), and Empty Maximal Space (EMS). These schemes are all related to boundary points of packed items $b\in\textbf{B}_{t}$ along the $d$ axis. We combine these boundary points to get the superset, namely Event Point (EV). See Appendix B for details and learning curves. We incorporate these schemes into our PCT model. Although the number of generated leaf nodes is reasonable, we only randomly intercept a subset $\textbf{L}_{sub_{t}}\subset\textbf{L}_{t}$ if $|\textbf{L}_{t}|$ exceeds a certain length, for saving computing resources. This interception length is constant during training and determined by a grid search (GS) during the test. See Appendix A for details. The performance comparisons are summarized in Table 2.
Table 2: Performance comparisons. “Uti.” and “Num.” represent the average space utilization and the average number of packed items, respectively. “Var.” (×$\times 10^{-3}$) refers to the variance of “Uti.” and “Gap” indicates the difference relative to the best “Uti.” across all methods. “Random” refers to placements selected randomly from all possible coordinates. DBL, LSAH, MACS, BR, and HM are heuristic methods proposed by Karabulut and Inceoglu (2004), Hu et al. (2017), Hu et al. (2020), Zhao et al. (2021), and Wang and Hauser (2019b). CDRL presents the constrained deep reinforcement learning method proposed by Zhao et al. (2022b).
[…]
@@
[…]
@@ 4.2 Benefits of Tree Presentation
[…]
@@
[…]
@@ 4.3 Performance on Continuous Dataset
[…]
@@
[…]
@@ 4.4 More Complex Practical Constraints
To further demonstrate that PCT effectively handles complex constraints, we conduct experiments extending PCT to online 3D-BPP with practical constraints, including isle friendliness, load balancing, and load bearing constraints proposed by Gzara et al. (2020), kinematic constraints (Martello et al., 2007), bridging constraints (Shin et al., 2016), and height uniformity:
Table 5: Online 3D-BPP with practical constraints. “Obj.” means the task-specific objective score, which is the primary focus of this comparison. Whether a higher or lower value is preferable depends on the task, with the preference indicated by arrows next to the task name. For load balancing and height uniformity tasks, the objective score is scaled by ×$\times 10^{3}$ and ×$\times 10^{2}$, respectively. “Imp.” indicates the percentage improvement in “Obj.” compared to a random placement policy.
@@
Obj. Imp. $\uparrow$ Uti. $\uparrow$ Num. $\uparrow$ Obj. Imp.$\uparrow$ Uti. $\uparrow$ Num. $\uparrow$ Obj. Imp.$\uparrow$ Uti. $\uparrow$ Num. $\uparrow$
@@
Kinematic Constraints $\uparrow$ CDRL $0.77$ $35.1\%$ $54.6\%$ $21.2$ $0.53$ $-31.2\%$ $60.0\%$ 23.1 $0.79$ $17.9\%$ $56.5\%$ $22.0$
@@
PCT 1.18 11.3% 69.2% 26.9 1.30 27.5% 80.6% 31.3 1.19 13.3% 69.0% 26.9
$\bullet$ Isle Friendliness stipulates that items of the same category should be packed closely as possible. The item weight is defined as $w_{t}=max(0,v_{t}-c\cdot dist(n_{t},\textbf{B}_{t}))$. Here $c$ is a constant and the objective function $dist(n_{t},\textbf{B}_{t})$ means the average distance between $n_{t}$ and the items of the same category in $\textbf{B}_{t}$. Category information is appended to the descriptors of $\textbf{B}_{t}$ and $n_{t}$. Four item categories are tested here.
$\bullet$ Load Balancing dictates that the packed items should have an even mass distribution within the bin. The item weight is set as $w_{t}=max(0,v_{t}-c\cdot var(n_{t},\textbf{B}_{t}))$. Object $var(n_{t},\textbf{B}_{t})$ is the variance of the mass distribution of the packed items on the bottom of the bin.
$\bullet$ Load-bearing Constraint considers that items placed above do not exert excessive weight on those below. The force on each item is simulated using physics engine (Coumans and Bai, 2016), and the item weight is $w_{t}=\max(0,v_{t}-c\cdot\text{f}(n_{t},\textbf{B}_{t}))$, where $\text{f}(n_{t},\textbf{B}_{t})$ represents the force born by $n_{t}$.
$\bullet$ Kinematic Constraint minimizes the impact of placed items on the robot’s subsequent motions. Instead of time-consuming motion planning (Görner et al., 2019), we use the safe position reward $V_{\text{safe}}$ proposed by Zhao et al. (2022b). The item weight is $w_{t}=v_{t}+c\cdot V_{\text{safe}}$.
$\bullet$ Bridging Constraint requires items to be stacked interlockingly, improving stability by distributing the center of gravity, increasing the contact area, and enhancing friction (Page, 1981). The number of items $b\in\textbf{B}_{t}$ that contribute to bridging $n_{t}$ is summed up as $\text{bridge}(n_{t},\textbf{B}_{t})$, and the item weight is $w_{t}=v_{t}+c\cdot\text{bridge}(n_{t},\textbf{B}_{t})$.
$\bullet$ Height Uniformity ensures even height distribution of items in the bin. The item weight is $w_{t}=\max(0,v_{t}-c\cdot H_{\text{var}})$, where $H_{\text{var}}$ is the variance of the heightmap inside the bin after packing the items $n_{t}$.
We adopt the learning-based method CDRL in Section 4.1 as the baseline, as heuristic methods primarily aim to maximize space utilization and are not flexible enough to handle additional constraints. The results are summarized in Table 5. PCT demonstrates strong adaptability to multiple complex constraints, achieving consistently higher objective scores. Its improvement over random placement (Imp.) clearly outperforms the baseline method, demonstrating that PCT effectively captures complex task constraints and is suitable for solving online packing in practical applications.

### v4

@@ 3.2 Markov Decision Process Formulation
The transition $\mathcal{P}(s_{t+1}|s_{t})$ is jointly determined by the current policy $\pi$ and the probability distribution of sampling items. Our online sequences are generated on the fly from an item set $\mathcal{I}$ in a uniform distribution. The transferability of our method on item sampling distributions different from the training one is discussed in Appendix D.
@@
Our reward function $R$ is defined as $r_{t}=c_{r}\cdot w_{t}$ once $n_{t}$ is inserted into PCT as an internal node successfully; otherwise, $r_{t}=0$ and the packing episode ends. Here, $c_{r}$ is a constant and $w_{t}$ is the weight of $n_{t}$. The choice of $w_{t}$ depends on the customized needs. For simplicity and clarity, unless otherwise noted, we set $w_{t}$ as the volume $v_{t}$ of $n_{t}$, where $v_{t}=s^{x}_{n}\cdot s^{y}_{n}\cdot s^{z}_{n}$.
@@ 3.3 Recursive Packing for Large-Scale BPP
Figure 3: Problem decomposition based on sliding windows, illustrated in the xozxoz plane. Low-resolution decomposition (a) results in the loss of solutions, while fine-grained partitioning (b) clearly increases computational overhead.
For problem decomposition, an intuitive approach is to divide the bin $C$ into uniform sections with resolution $r$ along each dimension, and maintain a sliding window that traverses the entire bin space in a convolution-like manner, resulting in sub-bins $\mathbf{c}=\{c^{1},...,c^{n}\}$. Each $c^{i}$, along with its overlapped packed items $\mathbf{B}^{i}$ and empty space $\mathbf{L}^{i}$, is treated as a sub-problem $\mathcal{T}^{i}$. While this decomposition is intuitive, it’s not aware of the in-bin spatial distribution, leading to potential degradation in solution quality. Figure 3 (a) gives a demonstration for this, solutions that exist in the original bin $C$ are no longer feasible in any of the decomposed sub-bin $c^{i}$. Although finer-grained partitioning can, to some extent, avoid this (Figure 3 (b)), it also leads to an $O(r^{6})$ space complexity, significantly increasing the number of sub-problems and computational costs. Even more, for problems requiring decisions in continuous domains, fine-grained partitioning is inherently infeasible.
@@
[…]
Figure 5: From the view of sub-bin $c_{1}$ (a), the current item is optimally placed, but for the global view of bin $C$ (b), unused space beneath the item exists, which can no longer be utilized.
@@
$l^{*}=\arg\max_{l\in\textbf{L}}\min_{c_{i}\in\mathbf{c}}\Phi(l,c_{i}).\vskip-4.0pt$ (9)
@@ 3.4 Uniform Planning for BPP Variations
Figure 7: Mainstream BPP variations–offline (a), buffering (b), and lookahead (c)–can be modeled as conflict-free search trees. We solve them via a unified framework (d) powered by a pre-trained PCT model with no additional adaptation.
@@
We first formalize offline 3D-BPP, where all items are selectable, as a planning problem. An intuitive approach is to perform a traversal tree search over all packing orders and positions for all items $\mathcal{I}$. Each path of the search tree represents a possible solution, and we choose the one with the highest accumulated space utilization $\sum_{i=0}^{|\mathcal{I}|}v_{i}$ for execution. This brute force search has a computational complexity of $O(|\mathcal{I}|!\cdot|\mathcal{A}|^{|\mathcal{I}|})$, where $|\mathcal{A}|$ represents action space size. Leveraging the pre-trained policy model $\pi_{\theta}$, which determines item position, the search can be simplified to only consider item order, as exhibited in Figure 7 (a), lowering the complexity to $O(|\mathcal{I}|!)$. The constraints for planning selectable items are as follows:
@@
    Items are placed based on the node sequence of the planned path. The planning is only conducted once.
Buffering packing can be considered a direct extension of offline packing, where unknown items should be additionally considered. These items cannot be explicitly included as tree nodes. We use the state value function $V(\cdot)$ to implicitly estimate their distribution and future values. For $s$ items within $Fov_{c}$, their placement order can still be enumerated during the planning, while items outside of $Fov_{c}$ are all modeled as a single leaf node in the path end, as shown in Figure 7 (b). The value of each path is calculated by the sum of item volumes in the path and the value of the leaf node, i.e., $\sum_{i=0}^{s}v_{i}+V(\cdot)$. Since future arrivals exist, the planning follows the principle of MBP, where only the first node of the selected path is executed, and the search reinitializes when a new item arrives. The additional planning constraints when introducing unknown items are summarized as:
@@
We can find that these planning constraints for different operation attributes are compatible and can be integrated into a unified framework, which can be applied to general packing scenarios where $p\geq 0$, $s\geq 1$, and $u\geq 0$, as illustrated in Figure 7 (d). We denote this unified search tree as Tree of Packing (ToP). To ensure that the planning meets the real-time requirements of industrial packing, we introduce Monte Carlo Tree Search (MCTS) (Silver et al., 2016) which reduces the time complexity of the brute-force search from $O(|p+s|!)$ to $O((p+s)\cdot m)$. Here $m$ is the number of sampled paths. During planning, MCTS at adjacent time steps may share the same part of paths (i.e., item sequences). We maintain a global cache that stores previously visited paths to avoid redundant computations. This approach nearly halves the decision time costs.
@@ 4 Experiments
In this section, we first present PCT performance combined with different leaf node expansion schemes. We then highlight the advantages of the structured packing representation, including improved node spatial relation representations and a more flexible action space. Next, we validate the effectiveness of PCT-driven planners in solving industrial packing problems, specifically large-scale packing and different variations of the BPP setting. Finally, we introduce our real-world packing robot in an industrial warehouse, carefully designed to meet constrained placement and transportation stability.
@@
Although there are very few online packing implementations publicly available, we still do our best to collect or reproduce various online 3D-BPP algorithms, both heuristic and learning-based, from potentially relevant literature. We help the heuristics to make pre-judgments of placement constraints, e.g., stability, in case of premature downtime. The learning-based agents are trained until there are no significant performance gains. All methods are implemented in Python and tested on 2000 instances with a desktop computer equipped with a Gold 5117 CPU and a GeForce TITAN V GPU. We evaluate performance based on the average space utilization and the average number of packed items, denoted as “Uti.” and “Num.” , respectively. We also report the variance ($\times 10^{-3}$) of “Uti.”, denoted as “Var.”. We include the “Gap” metric, which indicates the difference relative to the best “Uti.” across all methods.
@@
Some baselines (Karabulut and Inceoglu, 2004; Wang and Hauser, 2019b) need to traverse the entire coordinate space to find the optimal solution, and the running costs explode as the spatial discretization accuracy increases. To ensure that all algorithms are runnable within a reasonable period, we use the discrete dataset proposed by Zhao et al. (2021) without special declaration. The bin sizes $S^{d}$ are set to 10 with $d\in\{x,y,z\}$, and the item sizes $s^{d}\in\mathbb{Z}^{+}$ are not greater than $S^{d}/2$ to avoid over-simplification. Our performance on the continuous dataset will be reported in Section 4.3. Considering that there are many practical scenarios of 3D-BPP, we choose three representative ones:
Setting 1Setting\,1: Following Zhao et al. (2022b), stability of $\textbf{B}_{t}$ is verified when $n_{t}$ is placed. For robot manipulation convenience, only two horizontal orientations ($|\textbf{O}|=2$) are allowed for top-down placement.
Setting 2Setting\,2: Following Martello et al. (2000), item $n_{t}$ only needs to satisfy Constraints (2) and (3). Arbitrary orientation ($|\textbf{O}|=6$) is allowed here. This is the most common setting in the 3D-BPP literature.
Setting 3Setting\,3: Building on setting 1setting\,1, each item $n_{t}$ is assigned an additional density property $\rho$ uniformly sampled from $(0,1]$. This density information is incorporated into the descriptors of both $\textbf{B}_{t}$ and $n_{t}$.
@@ 4.1 Performance of PCT Policies
We first report the performance of PCT combined with different leaf node expansion schemes. Three existing schemes, which have proven to be both efficient and effective, are adopted here: Corner Point (CP), Extreme Point (EP), and Empty Maximal Space (EMS). These schemes are all related to boundary points of packed items $b\in\textbf{B}_{t}$ along the $d$ axis. We combine these boundary points to get the superset, namely Event Point (EV). See Appendix B for details and learning curves. We incorporate these schemes into our PCT model. Although the number of generated leaf nodes is reasonable, we only randomly intercept a subset $\textbf{L}_{sub_{t}}\subset\textbf{L}_{t}$ if $|\textbf{L}_{t}|$ exceeds a certain length, for saving computing resources. This interception length is constant during training and determined by a grid search (GS) during the test. See Appendix A for details. The performance comparisons are summarized in Table 2.
Table 2: Performance comparisons. “Random” refers to placements selected randomly from all coordinates. DBL, LSAH, MACS, BR, and HM are heuristics proposed by Karabulut and Inceoglu (2004), Hu et al. (2017), Hu et al. (2020), Zhao et al. (2021), and Wang and Hauser (2019b). CDRL presents the constrained deep reinforcement learning method proposed by Zhao et al. (2022b).
Method Setting 1Setting\,1 Setting 2Setting\,2 Setting 3Setting\,3
@@
The performance of Zhao et al. (2022b) deteriorates quickly in setting 2setting\,2 and setting 3setting\,3 due to the multiplying orientation space and insufficient state representation separately. Running costs, scalability performance, behavior understanding, and visualized results can be found in Appendix D. We also repeat the same experiment as Zhang et al. (2021), which packs items sampled from a pre-defined item set $|\mathcal{I}|$ = 64 in setting 2setting\,2. While the method of Zhang et al. (2021) packs on average $15.6$ items and achieves $67.0\%$ space utilization, our method packs $19.7$ items with a space utilization of $83.0\%$. Although EV sometimes yields better performance, we provide a detailed explanation in Appendix B that its computational complexity is quadratic, whereas EMS’s complexity is linear to internal nodes $|\textbf{B}|$. Therefore, we choose EMS as the default scheme. We provide a theoretical analysis of PCT in Appendix C: it selects candidates that are locally optimal in a tightness measure, while the DRL agent further determines the global decision, improving learning efficiency and performance.
@@ 4.2 Benefits of Tree Presentation
Presentation Setting 1Setting\,1 Setting 2Setting\,2 Setting 3Setting\,3
@@
If we ignore the spatial relations between the PCT nodes or only treat the state input as a flattened sequence, the performance of the learned policies will be severely degraded. The presence of B functions more on setting 1setting\,1 and setting 3setting\,3 since setting 2setting\,2 allows items to be packed in any empty spaces without considering constraints with internal nodes. This also confirms that a complete PCT representation is essential for online 3D-BPP of practical needs.
@@ 4.3 Performance on Continuous Dataset
where $\textbf{B}_{sub1}\neq\textbf{B}_{sub2},\textbf{B}_{sub1},\textbf{B}_{sub2}\in\textbf{B}_{t}$. This means the partition problem (Korf, 1998) has no solutions, and any packed items cannot form a new plane for providing support in the direction of gravity, leading to the available packing areas shrinking. Excessive diversity of $s^{z}$ will degenerate 3D-BPP into 1D-BPP as presented in the above toy demo. To prevent this degradation from leading to the underutilization of the bins, we sample $s^{z}$ from a finite set $\{0.1,0.2,\dots,0.5\}$ on setting 1setting\,1 and setting 3setting\,3 in this section.
@@
Method Setting 1Setting\,1 Setting 2Setting\,2 Setting 3Setting\,3
@@ 4.4 More Complex Practical Constraints
Table 5: Online 3D-BPP with practical constraints. “Obj.” means the task-specific objective score, which is the primary focus of this comparison. Whether a higher or lower value is preferable depends on the task, with the preference indicated by arrows next to the task name. For load balancing and height uniformity tasks, the objective score is scaled by $\times 10^{3}$ and $\times 10^{2}$, respectively. “Imp.” indicates the percentage improvement in “Obj.” compared to a random placement policy.
Constraints Method Setting 1Setting\,1 Setting 2Setting\,2 Setting 3Setting\,3
[…]
@@
Kinematic Constraints $\uparrow$ CDRL $0.77$ $35.1\%$ $54.6\%$ $21.2$ $0.53$ $31.2\%$ $60.0\%$ 23.1 $0.79$ $17.9\%$ $56.5\%$ $22.0$
@@
Bridging Constraints $\uparrow$ CDRL $1.09$ $2.8\%$ $59.3\%$ $23.0$ $1.03$ $1.0\%$ $60.8\%$ $23.4$ $1.07$ $1.9\%$ $59.7\%$ $23.1$
PCT 1.18 11.3% 69.2% 26.9 1.30 27.5% 80.6% 31.3 1.19 13.3% 69.0% 26.9
$\bullet$ Isle Friendliness stipulates that items of the same category should be packed as closely as possible. The constraint reward is $f=-\text{dist}(n_{t},\textbf{B}_{t})$, where $\text{dist}(n_{t},\textbf{B}_{t})$ measures the average distance between $n_{t}$ and the items of the same category in $\textbf{B}_{t}$. Category information is appended to the descriptors of $\textbf{B}_{t}$ and $n_{t}$. Four item categories are tested here.
$\bullet$ Load Balancing dictates that the packed items should have an even mass distribution. The constraint is $f=-\text{var}(n_{t},\textbf{B}_{t})$, where $\text{var}(n_{t},\textbf{B}_{t})$ is the variance of the mass distribution of the packed items on the bottom of the bin.
$\bullet$ Load-bearing Constraint considers that items placed above do not exert excessive weight on those below. The force $\text{bear}(b,\textbf{B}_{t})$ on each item $b$ is simulated using physics engine (Coumans and Bai, 2016). The constraint term is $f=-\mathbb{E}_{b\in{\textbf{B}_{t}}}\text{bear}(b,\textbf{B}_{t})$.
$\bullet$ Kinematic Constraint minimizes the impact of placed items on the robot’s subsequent motions. Instead of time-consuming motion planning (Görner et al., 2019), we use the safe position reward $V_{\text{safe}}$ (Zhao et al., 2022b) as constraint, with $f=V_{\text{safe}}$.
$\bullet$ Bridging Constraint requires packing items in a staggered manner to form bridges, improving stability by distributing the center of gravity (Page, 1981). The number of items $b\in\textbf{B}_{t}$ that contribute to bridging $n_{t}$ is summed up as $\text{bridge}(n_{t},\textbf{B}_{t})$, and the constraint reward is $f=\text{bridge}(n_{t},\textbf{B}_{t})$.
$\bullet$ Height Uniformity ensures even height distribution of items in the bin. The constraint reward is $f=-H_{\text{var}}$, where $H_{\text{var}}$ is the variance of the heightmap after packing $n_{t}$.